# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

### v4

@@ 4.6 Frequency of different data utilization
In the "Fuel Conditions" category, the NDVI is the most frequently utilized variable, with a frequency of 27, underscoring NDVI’s critical role in assessing vegetation health and fuel availability. Other significant variables include Tree Species, which appears 8 times, and LAI and Vegetation Type, each with a frequency of 6. These variables are essential for understanding the characteristics of biomass that contribute to wildfire risk. The presence of numerous variables with lower frequencies, such as Enhanced Vegetation Index (EVI) and NDWI, each used only once, highlights the complexity of this category. This diversity of variables reflects the multifaceted nature of fuel dynamics, requiring a broad range of factors to be considered in fire risk modeling, as depicted in Fig. 3(a).
@@
In the "Climate and Weather Conditions" category, Temperature emerges as the dominant variable, with a frequency of 65. Precipitation follows closely with 53 occurrences, and Wind is also significantly utilized, with a frequency of 42. These variables are fundamental in understanding fire behavior, particularly in their influence on fire ignition, spread, and intensity. Humidity, used 27 times, and Moisture, used 16 times, further contribute to this category, reflecting the central role of atmospheric conditions in fire risk assessments. Other variables, such as Fuel Moisture Code, utilized 11 times, and Vapor Pressure, with 9 occurrences, are less frequently employed but remain crucial in specific contexts. The data presented in Fig. 3(b) emphasizes that climate variables, particularly Temperature and Precipitation, are among the most critical factors in fire dynamics, with a wide array of related variables also contributing to detailed climatic modeling.
@@
Land Use and Land Cover (LULC) stands out as the most frequently utilized socio-economic variable, appearing 32 times, underscoring its importance in assessing how human activities and changes in land use influence fire risk. Road Proximity, with 23 occurrences, and Settlement Proximity, with 18, follow, indicating the significance of infrastructure and human settlement patterns in fire studies. Population Density, appearing 16 times, and Road Density, with a frequency of 4, also play roles, although to a lesser extent. The data in Fig. 3(c) suggest that while variables like Gross Domestic Product (GDP), used 5 times, and Distance to Cropland, also used 5 times, are relevant, they are more context-specific. This analysis indicates that socio-economic factors are essential for understanding the human impact on fire dynamics, particularly in areas where land use and human infrastructure intersect with natural landscapes.
@@
In the "Terrain and Hydrological Features" category, Slope is the most frequently used variable, appearing 46 times. This reflects the critical role of slope in influencing fire spread and intensity. Aspect, with a frequency of 34, and Elevation, with 32, are also key factors in understanding how topography affects fire behavior. These variables are particularly important in regions with complex terrain, where physical landscape features significantly impact fire dynamics. Plan Curvature, used 11 times, and TWI, used 10 times, are less frequently applied but remain relevant in specific contexts, often related to hydrological modeling. Other variables like Latitude, used 5 times, and Landform, with 3 occurrences, indicate their more specialized application. The analysis of these features highlights that while terrain and hydrological factors are crucial for understanding the physical environment’s impact on fires, certain variables like Slope, Aspect, and Elevation consistently prove to be more influential.
@@ 4.7 Open-source wildfire spread prediction datasets
The datasets presented in Table.LABEL:oa_datasets_description, including FireCube (Prapas et al., 2022; Kondylatos et al., 2023), Next Day Wildfire Spread (Huot et al., 2022), WildfireDB (Singla et al., 2020), WildfireSpreadTS (Gerard et al., 2023), CFSDS (Barber et al., 2024), and SeasFire Cube (Alonso et al., 2023), encompass various factors critical to wildfire spread prediction. These factors include spatial-temporal coverage of fuel conditions, climate and weather conditions, socio-economic factors, terrain, and historical records. Each dataset differs in the spatial-temporal coverage, and resolution of these factors, providing diverse perspectives and data for wildfire modeling.
First, the datasets vary significantly in their spatial and temporal resolutions. For instance, FireCube and Next Day Wildfire Spread focus on coarse-resolution data with a spatial coverage of 1 km for daily observations. WildfireDB offers more localized coverage at 375 m but over a broader temporal range (2012-2018). The CFSDS dataset provides highly detailed data at 180m resolution, though its temporal span is longer (2002-2021). SeasFire Cube offers the broadest coverage globally, albeit at a coarser resolution of 0.25 degrees and 8-day intervals.
@@
It is worth noting that, as shown in Table.LABEL:oa_datasets_description, the number of available open-source wildfire spread datasets is limited, with most coverage concentrated in the United States and Mediterranean, while other regions, such as the wildfire-prone Canada and Australia, are underrepresented. Additionally, some studies using these datasets have overly simplistic considerations of fuel conditions, such as relying solely on vegetation indices as proxies. This approach fails to accurately and comprehensively describe the state of fuel load, moisture content, and continuity. Similarly, using MODIS or VIIRS to record historical burned areas lacks precision, potentially overlooking small fire spots, offering low spatial and temporal resolution, and being prone to false alarms (Zubkova et al., 2024). Furthermore, there is a lack of quantitative evaluation of how the accuracy of different products impacts the accuracy of wildfire spread predictions.
@@ 5 Features Collinearity and Attribution
The relationships between wildfires and driving environmental factors have always been a focal point in wildfire prediction research. Considering the interplay between wildfire occurrences and factors such as fuel conditions, climate, meteorological variables, topographical features, and human activities is crucial. However, including highly collinear independent variables in a model may impair its performance (Chen et al., 2019). Moreover, when two or more independent variables are significantly correlated, they provide redundant information in regression analysis, complicating the interpretation of each variable’s individual impact on the dependent variable. Therefore, it is essential to examine the collinearity and contributions of different features to provide predictive insights for wildfire management.
@@ 5.1 Assessment of Feature Collinearity
Common metrics for assessing data collinearity in wildfire risk prediction include the Variance Inflation Factor (VIF) and tolerance (Hair Jr et al., 1986; Liao and Valliant, 2012a). VIF quantifies how much the variance of a regression coefficient is inflated due to collinearity, while tolerance represents the proportion of variance in a predictor that is not explained by other predictors in the model. The formulas are as follows:
@@
For example, Nami et al. (2018) conducted a collinearity analysis using VIF and tolerance in their study on wildfire risk mapping in the northern Hyrcanian ecoregion of Iran. They found significant collinearity between the TWI and slope with other parameters, leading to their exclusion from subsequent modeling efforts.
Similarly, Jaafari et al. (2019) assessed the collinearity of various predictors using VIF and tolerance in their study in the western Zagros ecoregion of Iran. They discovered that the VIF values for TWI and slope were below 5, and tolerance values were above 0.2, indicating insignificant collinearity with other predictors (Liao and Valliant, 2012b). Hong et al. (2019), in their modeling of wildfire susceptibility in Huichang County, China, used VIF and tolerance to assess the collinearity of multiple predictors and found no significant collinearity between slope and other features.
@@
In summary, different studies have shown varying levels of collinearity between features, indicating that no consistent conclusion can be drawn regarding the relationships between specific features across different wildfire risk prediction models.
@@ 5.2 Feature attribution
Firstly, PFI quantitatively analyzes feature importance by randomly permuting each feature and evaluating the change in model performance. For example, Rubí and Gondim (2023) used PFI to assess the contributions of meteorological, fuel, terrain, and socioeconomic factors in models such as ANN, SVM, NB, KNN, LR, LogLR, and AdaBoost in their study in the Federal District of Brazil. They found that analyzing individual variables might lead to contradictory results. For instance, NDVI was more important for models like AdaBoost and ANN, whereas its influence was relatively low or nonexistent in other models such as LR and SVM. This finding highlights the necessity of considering multiple variables in predictive modeling and the varying importance of different models. Additionally, Shadrin et al. (2024a) evaluated the impact of specific feature values on the final test results by setting them to zero during model testing. Similarly, Chen et al. (2024b) calculated the importance of each feature by measuring the variation in prediction error when feature values were perturbed.
@@
6 Fire danger rating systems
Wildfire hazards are characterized by strong nonlinearity and randomness, making it challenging to calculate wildfire risk using simple statistical models. At the same time, developing accurate physical models to precisely estimate the relationship between wildfire risk and factors such as fuel, weather, terrain, and human activities is also difficult. Consequently, adaptive and scientifically grounded wildfire rating systems have been proposed and have remained a critical component of global fire management for over a century (Hardy and Hardy, 2007). These systems have evolved from early dryness indices into complex models that predict spatial and temporal variations in wildfire ignition probability, spread rate, and heat release/fire intensity by characterizing the interactions among fuel, weather, and terrain. This evolution has provided crucial decision support for wildfire management at local, regional, national, and sometimes even international scales (Jolly et al., 2024). According to Zacharakis and Tsihrintzis (2023a), commonly used wildfire danger rating models today include systems developed in Canada, the United States, and Australia.
Figure 4: Illustration of FWI system.
For example, Canada developed a meteorology-based fire danger rating method in 1968, known as the Canadian Forest Fire Danger Rating System (CFFDRS). This system comprises four main subsystems: the FWI, Fire Behavior Prediction (FBP), Fire Occurrence Prediction (FOP), and Auxiliary Fuel Moisture (AFM) (Van Wagner et al., 1974; Stocks et al., 1989b; Lawson and Armitage, 2008). Among these, the FWI system is the most widely used (Fig. 4). The FWI system, based on the theory of time lag and equilibrium moisture content, calculates changes in fuel moisture codes (i.e., Fine Fuel Moisture Code (FFMC), Duff Moisture Code (DMC), and Drought Code (DC)) and fire behavior indices (i.e., Initial Spread Index (ISI) and Buildup Index (BI)) according to weather conditions. These indices are then used to classify potential wildfire danger levels in forests based on the moisture content of fuels at different locations or sizes (Stocks et al., 1989a).
Furthermore, the FWI system has been proven applicable in other countries or regions (Alexander, 1989; Fogarty et al., 1998; López et al., 2002; Taylor and Alexander, 2006; Dimitrakopoulos et al., 2011; Giannakopoulos et al., 2012). For instance, (Viegas et al., 1999) demonstrated that the FWI system is suitable for Europe, and it has been widely adopted by European countries as the optimal method for assessing wildfire danger (San-Miguel-Ayanz et al., 2018). Some countries have also developed their wildfire rating systems based on the FWI system. For example, South Africa developed the Lowveld Fire Danger Index (LFDI) based on the FWI, as shown in Eq. (4). Research has proven that this index is also applicable to the Mediterranean region of Greece (Cavalcante et al., 2021). New Zealand’s wildfire rating system incorporates the Rate of Spread (ROS) and Head Fire Intensity (HFI) into the FWI system (Gregor and David, 2017). Similarly, Groot et al. (2007) developed an early wildfire risk warning system for Indonesia and Malaysia based on the FWI and FBP subsystems of the CFFDRS. This system includes a smoke potential indicator based on the DC, an ignition potential indicator based on the FFMC, and a difficulty of control indicator for grassland fires based on the ISI of the FWI System. The ISI-based indicator was developed using the grass fuel model of the FBP System, along with a standard grass fuel load and curing level estimated from previous Indonesian studies.
In comparison, the United States National Fire Danger Rating System (USNFDRS) (Deeming, 1972) evaluates fire risk by considering not only meteorological factors, but also various terrain types, fuel types, and fuel models. By inputting meteorological data, fuel moisture, and terrain into the model, four output components are obtained: the Spread Component (SC), Energy Release Component (ERC), Burning Index, and Ignition Component (IC). The first three indices, based on combustion physics, correspond to fire behavior characteristics, while the IC provides an estimate of fire danger levels (Zacharakis and Tsihrintzis, 2023a). The system was revised in 1978 and 1988 to address issues such as poor response to prolonged droughts and the inability to capture real-time fuel dynamics (Deeming et al., 1977) and inadequate performance in moist environments (Burgan, 1988). Recently, Jolly et al. (2024) updated the USNFDRS introduced in 1988 by emphasizing the provision of more detailed, fuel-type-dependent combustion condition assessments, proposing a new fuel moisture model, and simplifying the fuel model.
Figure 5: Illustration of AFDRS system (Hollis et al., 2024).
In Australia, widely used fire danger rating systems include the early McArthur’s Forest Fire Danger Model (McArthur, 1966, 1967, 1973; Noble et al., 1980) and the Australia Fire Danger Rating System (AFDRS) introduced in 2022. The first system derived fire danger ratings based on predicted wildfire spread rates and the difficulty of suppressing fires burning in grasslands (McArthur, 1960) and dry sclerophyll forests (McArthur, 1958). This index, originally designed for eastern Australia in the 1950s and 1960s, underwent several revisions, culminating in the final Mark 5 version. Depending on vegetation type, it is divided into the Forest Fire Danger Index (FFDI) and the Grassland Fire Danger Index (GFDI) (Zacharakis and Tsihrintzis, 2023a). However, Hollis et al. (2024) pointed out that McArthur’s Forest Fire Danger Model could not utilize contemporary knowledge of fire behavior, such as new wildfire spread models (Cheney et al., 2012; Cruz et al., 2013; Anderson et al., 2015) and fire-atmosphere coupling (Mills and McCaw, 2010), and that its application has extended beyond its original design (Fogarty et al., 2010). Additionally, this system, with its low spatial resolution, cannot provide specific location-based wildfire risk recommendations. Finally, FFDI and GFDI cannot describe wildfire risks for vegetation types like shrublands (Anderson et al., 2015) and Spinifex (Burrows et al., 2018). Therefore, scientists and representatives from fire and land management agencies across all Australian jurisdictions collaborated to develop the new AFDRS. This system includes six factors: weather, fuel, fire behavior, ignition likelihood, suppression capability, and fire impact. It is designed for consistent and standardized use across Australia’s states and territories, eliminating the need for ad hoc local adjustments. The main components of this system are shown in Fig. (5), which determines fire danger by establishing a forward spread rate model (Cruz et al., 2015) and incorporating hourly forecast weather data on a 1.5 km grid across Australia, as well as fuel type and fuel state, to calculate forward spread rates, fireline intensity, and other relevant information.
@@ 7 Deep Learning Methods for Wildfire Risk Prediction
Traditional machine learning methods often use shallow architectures that fail to exploit spatial patterns effectively, limiting feature extraction and classification accuracy (Zhang et al., 2018). In contrast, deep learning models, despite their higher data requirements, can automatically extract complex features through hierarchical structures, reducing the need for manual feature engineering and enhancing generalizability. Their larger parameter capacity also allows them to manage large datasets and minimize overfitting risks. Although deep learning has become increasingly popular for wildfire risk prediction, reviews such as Jain et al. (2020) and Zacharakis and Tsihrintzis (2023b) have only partially covered models like Recurrent Neural Networks (RNNs) and CNNs without a comprehensive systematization.
@@
Deep learning models for spatio-temporal prediction often combine CNNs, Transformers, GCNs, and graph neural networks (GNNs) with RNNs, LSTMs, or GRUs, enhancing resolution with higher-resolution data over extended sequences. Despite advances in GPU technology supporting larger data scales, these models still struggle with efficiency and large spatiotemporal scales compared to newer models like Mamba (Gu and Dao, 2024). Additionally, many models use multi-branch networks that rely solely on images or nodes, missing the benefits of multimodal approaches like incorporating textual information.
While previous reviews (e.g., Jain et al. (2020); Zacharakis and Tsihrintzis (2023b)) have examined machine learning in wildfire science, they often omit post-processing techniques essential for converting model outputs into actionable risk information. Directly using these outputs can be misleading due to class imbalances and variability between models. This section also introduces methods for classifying and calibrating deep learning wildfire risk predictions to address these challenges.

### v5

@@ 4.6 Frequency of Different Data Utilization
In the "Fuel Conditions" category, the NDVI is the most frequently utilized variable, with a frequency of 27, underscoring NDVI’s critical role in assessing vegetation health and fuel availability. Other significant variables include Tree Species, which appears 8 times, and LAI and Vegetation Type, each with a frequency of 6. These variables are essential for understanding the characteristics of biomass that contribute to wildfire risk. The presence of numerous variables with lower frequencies, such as Enhanced Vegetation Index (EVI) and NDWI, each used only once, highlights the complexity of this category. This diversity of variables reflects the multifaceted nature of fuel dynamics, requiring a broad range of factors to be considered in fire risk modeling, as depicted in Fig. 3.
@@
In the "Climate and Weather Conditions" category, Temperature emerges as the dominant variable, with a frequency of 65. Precipitation follows closely with 53 occurrences, and Wind is also significantly utilized, with a frequency of 42. These variables are fundamental in understanding fire behavior, particularly in their influence on fire ignition, spread, and intensity. Humidity, used 27 times, and Moisture, used 16 times, further contribute to this category, reflecting the central role of atmospheric conditions in fire risk assessments. Other variables, such as Fuel Moisture Code, utilized 11 times, and Vapor Pressure, with 9 occurrences, are less frequently employed but remain crucial in specific contexts. The data presented in Fig. 3 emphasizes that climate variables, particularly Temperature and Precipitation, are among the most critical factors in fire dynamics, with a wide array of related variables also contributing to detailed climatic modeling.
@@
Land Use and Land Cover stands out as the most frequently utilized socio-economic variable, appearing 32 times, underscoring its importance in assessing how human activities and changes in land use influence fire risk. Road Proximity, with 23 occurrences, and Settlement Proximity, with 18, follow, indicating the significance of infrastructure and human settlement patterns in fire studies. Population Density, appearing 16 times, and Road Density, with a frequency of 4, also play roles, although to a lesser extent. The data in Fig. 3 suggest that while variables like Gross Domestic Product (GDP) and Distance to Cropland—each used 5 times—are relevant, they are more context-specific. This analysis indicates that socio-economic factors are essential for understanding the human impact on fire dynamics, particularly in areas where land use and human infrastructure intersect with natural landscapes.
@@
In the "Terrain and Hydrological Features" category, Slope is the most frequently used variable, appearing 46 times. This reflects the critical role of slope in influencing fire spread and intensity. Aspect, with a frequency of 34, and Elevation, with 32, are also key factors in understanding how topography affects fire behavior. These variables are particularly important in regions with complex terrain, where physical landscape features significantly impact fire dynamics. Plan Curvature, used 11 times, and TWI, used 10 times, are less frequently applied but remain relevant in specific contexts, often related to hydrological modeling. Other variables like Latitude, used 5 times, and Landform, with 3 occurrences, indicate their more specialized application. The analysis of these features highlights that, while terrain and hydrological factors are crucial for understanding the physical environment’s impact on fires, certain variables like Slope, Aspect, and Elevation consistently prove to be more influential.
@@ 4.7 Open-Source Wildfire Spread Prediction Datasets
The datasets presented in Table  LABEL:oa_datasets_description, including FireCube (Prapas et al., 2022; Kondylatos et al., 2023), Next Day Wildfire Spread (Huot et al., 2022), WildfireDB (Singla et al., 2020), WildfireSpreadTS (Gerard et al., 2023), CFSDS (Barber et al., 2024), and SeasFire Cube (Alonso et al., 2023), encompass various factors critical to wildfire spread prediction. These factors include spatial-temporal coverage of fuel conditions, climate and weather conditions, socio-economic factors, terrain, and historical records. Each dataset differs in the spatial-temporal coverage, and resolution of these factors, providing diverse perspectives and data for wildfire modeling.
First, the datasets vary significantly in their spatial and temporal resolutions. For instance, FireCube and Next Day Wildfire Spread focus on coarse-resolution data with a spatial resolution of 1 km for daily observations. WildfireDB offers more localized coverage at 375 m but over a broader temporal range (2012-2018). The CFSDS dataset provides highly detailed data at 180m resolution, though its temporal span is longer (2002-2021). SeasFire Cube offers the broadest coverage globally, albeit at a coarser resolution of 0.25 degrees and 8-day intervals.
@@
It is worth noting that, as shown in Table LABEL:oa_datasets_description, the number of available open-source wildfire spread datasets is limited, with most coverage concentrated in the United States and Mediterranean, while other regions, such as the wildfire-prone Canada and Australia, are underrepresented. Additionally, some studies using these datasets have overly simplistic considerations of fuel conditions, such as relying solely on vegetation indices as proxies. This approach fails to accurately and comprehensively describe the state of fuel load, moisture content, and continuity. Similarly, using MODIS or VIIRS to record historical burned areas lacks precision. These datasets may may overlook small fire spots, offer low spatial and temporal resolution, and be prone to false alarms (Zubkova et al., 2024). Furthermore, there is a lack of quantitative evaluation of how the accuracy of different products impacts the accuracy of wildfire spread predictions.
@@ 5 Features Collinearity and Attribution
The relationship between wildfire risk and its driving factors has long been a focal point in wildfire prediction research. Understanding the interactions between wildfire occurrence and factors such as fuel conditions, climate, meteorological variables, terrain characteristics, and human activities is critical for model design, data selection, and supporting wildfire management decisions (Aldersley et al., 2011). However, the inherent "black-box" nature of deep learning models poses significant challenges in interpreting the contribution of individual factors to wildfire risk predictions (Castelvecchi, 2016). This section systematically reviews commonly used feature attribution methods in deep learning-based wildfire risk prediction models, highlighting their potential and limitations.
Furthermore, wildfire risk prediction models often incorporate multiple driving factors, some of which may exhibit a high degree of multicollinearity. Highly correlated independent variables can introduce redundancy, complicating the isolation and interpretation of their individual contributions to the dependent variable (Chen et al., 2019). Systematically assessing multicollinearity and quantifying the individual contributions of these factors are essential for improving model reliability and providing actionable insights for effective wildfire management. Additionally, commonly used features in wildfire risk prediction models—such as precipitation, temperature, wind speed, terrain attributes, and vegetation indices—often overlap across studies. Investigating whether these features share inherent collinearity, independent of spatial and temporal distributions, is particularly valuable. This analysis could guide feature selection in future research, especially in data-constrained scenarios, ultimately enhancing the efficiency and accuracy of model development.
@@ 5.1 Assessment of Feature Collinearity
Multicollinearity refers to the presence of approximate linear dependencies among two or more predictor variables, meaning that one predictor can be linearly represented by a combination of others (Senawi et al., 2017). This phenomenon leads to several issues, including inflated standard errors of coefficients, coefficients with signs inconsistent with theoretical expectations, statistically insignificant parameters despite strong correlations between predictor variables and the outcome, and excessively large correlation coefficients relative to the explanatory power of the model (Chan et al., 2022).
Under multicollinearity, regression methods often produce redundant, unstable models with poor generalization performance, making it difficult to isolate the marginal effects of individual predictors (Katrutsa and Strijov, 2017; Chan et al., 2022). Although deep learning models—which do not rely on linear parameter estimation—are less affected in terms of predictive performance, multicollinearity can still adversely impact the accuracy and stability of interpretability in predictive analysis. This, in turn, reduces the overall efficiency of model interpretation and analysis (Salih et al., 2024).
To evaluate collinearity in wildfire risk prediction, common metrics such as the Variance Inflation Factor (VIF) and tolerance are widely employed (Hair Jr et al., 1986; Liao and Valliant, 2012). VIF measures the extent to which the variance of a regression coefficient is inflated due to collinearity, while tolerance quantifies the proportion of variance in a predictor that is not explained by other predictors in the model. These metrics are critical for identifying and mitigating collinearity-related issues, ensuring robust model development and interpretation. The formulas for these metrics are as follows:
@@
VIF had been widely used in wildfire risk predictions. For example, Nami et al. (2018) conducted a collinearity analysis using VIF and tolerance in their study on wildfire risk mapping in the northern Hyrcanian ecoregion of Iran. They found significant collinearity between the TWI and slope with other parameters, leading to their exclusion from subsequent modeling efforts. In contrast, Jaafari et al. (2019) assessed the collinearity of various predictors using VIF and tolerance in their study in the western Zagros ecoregion of Iran. They discovered that the VIF values for TWI and slope were below 5, and tolerance values were above 0.2, indicating insignificant collinearity with other predictors (Liao and Valliant, 2012). Hong et al. (2019), in their modeling of wildfire susceptibility in Huichang County, China, used VIF and tolerance to assess the collinearity of multiple predictors and found no significant collinearity between slope and other features.
@@
In summary, various studies have revealed differing levels of multicollinearity among features, highlighting the challenge of drawing consistent conclusions about the relationships between specific features in wildfire risk prediction models. Nevertheless, feature multicollinearity analysis remains critical as it facilitates dimensionality reduction, thereby accelerating model training and inference. More importantly, addressing multicollinearity can significantly improve the accuracy, reliability, and stability of feature attribution, particularly when employing attribution methods that assume feature independence (Salih et al., 2024).
@@ 5.2 Feature Attribution
Firstly, PFI quantitatively analyzes feature importance by randomly permuting each feature and evaluating the change in model performance. For example, Rubí and Gondim (2023) used PFI to assess the contributions of meteorological, fuel, terrain, and socioeconomic factors in models such as ANN, SVM, NB, KNN, LR, LogLR, and AdaBoost in their study in the Federal District of Brazil. They found that analyzing individual variables might lead to contradictory results. For instance, NDVI was more important for models like AdaBoost and ANN, whereas its influence was relatively low or nonexistent in other models such as LR and SVM. This finding highlights the necessity of considering multiple variables in predictive modeling and the varying importance of different models. Additionally, Shadrin et al. (2024) evaluated the impact of specific feature values on the final test results by setting them to zero during model testing. Similarly, Chen et al. (2024b) calculated the importance of each feature by measuring the variation in prediction error when feature values were perturbed.
@@ 6 Deep Learning Methods for Wildfire Risk Prediction
Traditional machine learning methods often use shallow architectures that fail to exploit spatial patterns effectively, limiting feature extraction and classification accuracy (Zhang et al., 2018). In contrast, deep learning models, despite their higher data requirements, can automatically extract complex features through hierarchical structures, reducing the need for manual feature engineering and enhancing generalizability. Their larger parameter capacity also allows them to manage large datasets and minimize overfitting risks. Although deep learning has become increasingly popular for wildfire risk prediction, reviews such as Jain et al. (2020) and Zacharakis and Tsihrintzis (2023) have only partially covered models like Recurrent Neural Networks (RNNs) and CNNs without a comprehensive systematization.
@@
Deep learning models for spatio-temporal prediction often combine CNNs, Transformers, GCNs, and graph neural networks (GNNs) with RNNs, LSTMs, or GRUs, enhancing resolution with higher-resolution data over extended sequences. Despite advances in GPU technology supporting larger data scales, these models still struggle with efficiency and large spatiotemporal scales compared to newer models like Mamba (Gu and Dao, 2023). Additionally, many models use multi-branch networks that rely solely on images or nodes, missing the benefits of multimodal approaches like incorporating textual information.
While previous reviews (e.g., Jain et al. (2020); Zacharakis and Tsihrintzis (2023)) have examined machine learning in wildfire science, they often omit post-processing techniques essential for converting model outputs into actionable risk information. Directly using these outputs can be misleading due to class imbalances and variability between models. This section also introduces methods for classifying and calibrating deep learning wildfire risk predictions to address these challenges.